Solve the system by addition.
–3x – 5y = –5
x – 2y = 9

Answer by funmath(2933)

Solve the system by addition.
–3x – 5y = –5
x – 2y = 9
Multiply the second equation by 3 and add it to the first equation to eliminate the x's and solve for y:
3(x-2y)=3(9)--->3x-6y=27
:
-3x-5y=-5
3x-6y=27
----------
0-11y=22
-11y=22
-11y/-11=22/-11
y=-2
Plug -2 in for y in either equation and solve for x, I'm going with the second equation because it's slightly easier.
x-2(-2)=9
x+4=9
x+4-4=9-4
x=5
The solution (x,y)=(5,-2)
Happy Calculating!!!!